Exhibit 14.4

The current stock price of ABC Corporation is $53.50. ABC Corporation has the following put and call option prices that expire six months from today. The risk-free rate of return is 5 percent, and the expected return on the market is 11 percent.

Exercise Price

Put Price

Call Price

50

$1.50

$5.75

55

$3.25

---

Refer to Exhibit 14.4. What should the price be of a call option that expires six months from today with an exercise price of $55?
Answer #1

Using put call parity
Call price C=S+P-X/(1+r)^t=53.50+3.25-55/(1+5%)^(6/12)=3.0754960